# StyleSinger 2: Zero-Shot Singing Voice Synthesis with Style Transfer and Multi-Level Style Control

### v1

@@ Abstract
Zero-shot singing voice synthesis (SVS) with style transfer and style control aims to generate high-quality singing voices with unseen timbres and styles (including singing method, emotion, rhythm, technique, and pronunciation) from audio and text prompts. However, the multifaceted nature of singing styles poses a significant challenge for effective modeling, transfer, and control. Furthermore, current SVS models often fail to generate singing voices rich in stylistic nuances for unseen singers. To address these challenges, we introduce StyleSinger 2, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. Specifically, StyleSinger 2 proposes three primary modules: 1) the clustering style encoder employs a clustering vector quantization model to stably condense style information into a compact latent space; 2) the Style and Duration Language Model (S&D-LM) concurrently predicts style information and phoneme duration, which benefits both; 3) the style adaptive decoder uses a novel mel-style adaptive normalization method to generate singing voices with enhanced details. Experimental results show that StyleSinger 2 outperforms all baseline models in synthesis quality, singer similarity, and style controllability across various tasks, including zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing style transfer. Singing voice samples can be accessed at https://stylesinger2.github.io/.
StyleSinger 2: Zero-Shot Singing Voice Synthesis with Style Transfer
@@
Yu Zhang Ziyue Jiang Ruiqi Li Changhao Pan Jinzheng He Rongjie HuangChuxin WangZhou Zhaothanks: Corresponding Author Zhejiang University {yuzhang34,ziyuejiang,zhaozhou}@zju.edu.cn
@@ 1 Introduction
To address these challenges, we introduce StyleSinger 2, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. StyleSinger 2 transfers and controls styles (like singing methods, emotion, rhythm, techniques, and pronunciation) from audio and text prompts to synthesize high-quality singing voices. To model diverse styles (like singing methods, emotion, rhythm, technique, and pronunciation), we propose the clustering style encoder, which uses a clustering vector quantization (CVQ) model to condense style information into a compact latent space, thus facilitating subsequent predictions, as well as enhance both training stability and reconstruction quality. For style transfer and control, we introduce the Style and Duration Language Model (S&D-LM). The S&D-LM incorporates a multi-task language module using audio and text prompts to concurrently predict style information and phoneme duration, thereby enhancing both. To generate singing voices rich in stylistic nuances, we introduce the style adaptive decoder, which employs a novel mel-style adaptive normalization method to refine mel-spectrograms with decoupled style information. Our experimental results show that StyleSinger 2 outperforms other baseline models in metrics including synthesis quality, singer similarity, and style controllability across various tasks, including zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing (STS) style transfer. Our main contributions can be summarized as follows:
  • We present StyleSinger 2, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. StyleSinger 2 excels in personalized and controllable SVS tasks.
@@
  • Experimental results show that StyleSinger 2 surpasses baseline models in synthesis quality, singer similarity, and style controllability across various tasks: zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing style transfer.
@@ 2.2 Style Modeling, Transfer and Control
3 StyleSinger 2
Figure 1: The architecture of StyleSinger 2. In Figure (a), S&D-LM represents the Style and Duration Language Model, and LR stands for length regulator. In Figure (b), the S&D-LM autoregressively predicts style information and phoneme duration. In Figure (c), intermediate mel-spectrograms are refined with style information in the style adaptive decoder. In Figure (d), the clustering style encoder extracts style information from mel-spectrograms.
In this section, we first overview the proposed StyleSinger 2. Then, we introduce several critical components, including the clustering style encoder, the style adaptive decoder, and the Style and Duration Language Model (S&D-LM). Finally, we elaborate on the training and inference procedures.
@@ 3.1 Overview
The architecture of StyleSinger 2 is depicted in Figure 1(a). We disentangle singing voices into separate representations for content, style (including singing method, emotion, rhythm, technique, and pronunciation), and timbre. For content representation, lyrics are encoded through a phoneme encoder, while a note encoder captures musical notes. For style representation, we use clustering vector quantization (CVQ) in the clustering style encoder to stably condense style information into a compact latent space, thus facilitating subsequent predictions. For timbre representation, we feed a prompt mel-spectrogram, sampled from different audio of the same singer, into the timbre encoder to obtain a one-dimensional timbre vector, disentangling timbre from other information. Then, we utilize the Style and Duration Language Model (S&D-LM) to predict style information and phoneme duration. Since styles and duration of singing voices are closely related, a composite module benefits both. Moreover, the S&D-LM achieves both style transfer and style control with audio and text prompts. Next, we employ the pitch diffusion predictor for F0 prediction and the style adaptive decoder to generate the target mel-spectrogram. The style adaptive decoder generates intricately detailed singing voices using a novel mel-style adaptive normalization method. During training, we train the clustering style encoder in the first phase and S&D-LM in the second phase. During inference, we can input audio prompts or text prompts to S&D-LM for style transfer or control. Please refer to Appendix A for more details.
@@ 3.5 Training and Inference Procedures
Training Procedures The final loss terms of StyleSinger 2 in the training phase consist of the following parts: 1) CVQ loss $\mathcal{L}_{CVQ}$: the CVQ loss for the clustering style encoder; 2) Pitch reconstruction loss $\mathcal{L}_{gdiff},\mathcal{L}_{mdiff}$: the Gaussian diffusion loss and the multinomial diffusion loss between the predicted and the GT pitch spectrogram for the pitch diffusion predictor; 3) Mel reconstruction loss $\mathcal{L}_{mae},\mathcal{L}_{ssim}$: the MAE loss and the SSIM loss between the predicted and the GT mel-spectrogram for the style adaptive decoder. 4) Duration prediction loss $\mathcal{L}_{dur}$: the MSE loss between the predicted and the GT phoneme-level duration in log scale for S&D-LM in the teacher-forcing mode; 5) Style prediction loss $\mathcal{L}_{style}$: the cross-entropy loss between the predicted and the GT style information for S&D-LM in the teacher-forcing mode.
Figure 2: Inference procedure of StyleSinger 2. In Figure (a), the S&D-LM extracts information from the audio prompt to predict the target style information and phoneme duration, while in Figure (b), the S&D-LM uses multi-level text prompts to predict them.
@@ 4.1 Experimental Setup
In this section, we present the datasets utilized by StyleSinger 2, delve into the implementation and training details, discuss the evaluation methodologies, and introduce the baseline models.
Dataset We use the open-source singing dataset with style annotations, GTSinger (Zhang et al., 2024b), specifically its Chinese and English subset (5 singers, 36 hours of Chinese and English singing and speech). We also enrich our data with M4Singer (Zhang et al., 2022a) (20 singers, 30 hours of Chinese singing), OpenSinger (Huang et al., 2021) (93 singers, 85 hours of Chinese singing), AISHELL-3 (Shi et al., 2021) (218 singers, 85 hours of Chinese speech), and a subset of PopBuTFy (Liu et al., 2022b) (20 singers, 18 hours of English speech and singing). Then, we manually annotate these singing data with global style labels (singing method and emotion) and six phoneme-level singing techniques. Subsequently, we randomly chose 40 singers as the unseen test set to evaluate StyleSinger 2 in the zero-shot scenario for all tasks. Notably, our dataset partitioning carefully ensures that both training and test sets for all tasks include cross-lingual singing and speech data. Please refer to Appendix B for more details.
@@
Evaluation Details We use both objective and subjective evaluation metrics to validate the performance of StyleSinger 2. For subjective metrics, we conduct the MOS (mean opinion score) and CMOS (comparative mean opinion score) evaluation. We employ the MOS-Q to judge synthesis quality (including clarity, naturalness, and rich stylistic details), MOS-S to assess singer similarity (in terms of timbre and styles) between the result and prompt, and MOS-C to evaluate style controllability (accuracy and expressiveness of style control). Both these metrics are rated from 1 to 5 and reported with 95% confidence intervals. In the ablation study, we employ CMOS-Q to gauge synthesis quality, and CMOS-S to evaluate singer similarity. For objective metrics, we use Singer Cosine Similarity (Cos) to judge singer similarity, and Mean Cepstral Distortion (MCD) along with F0 Frame Error (FFE) to quantify synthesis quality. Please refer to Appendix C for more details.
@@
Baseline Models We conduct a comprehensive comparative analysis of synthesis quality, style controllability, and singer similarity for StyleSinger 2 against several baseline models. Initially, we evaluate our model against the ground truth (GT) and the audio generated by HiFi-GAN (GT (vocoder)). Additionally, we examine StyleSinger 2 with two high-performing speech models that conduct style transfer: YourTTS (Casanova et al., 2022) and Mega-TTS (Jiang et al., 2023). To ensure a fair comparison for singing tasks, we enhance these models with a note encoder to process music scores and train them on speech and singing data. Subsequently, we also compare with the best traditional SVS model, RMSSinger (He et al., 2023). Furthermore, we assess StyleSinger 2’s performance against StyleSinger (Zhang et al., 2024a), the first model for zero-shot SVS with style transfer.
@@ 4.2 Main Results
Figure 3: Mel-spectrograms depicting the results of zero-shot style transfer. StyleSinger 2 effectively captures the rhythm and pronunciation in red boxes, along with the vibrato technique and rhythm in yellow boxes.
@@
Zero-Shot Style Transfer To assess the performance of StyleSinger 2 and baseline models in the zero-shot style transfer task, we randomly select samples with unseen singers from the test set as targets and different utterances from the same singers to form prompts. As shown in Table 1, we have the following findings: 1) StyleSinger 2 exhibits outstanding synthesis quality, as indicated by the highest MOS-Q and the lowest FFE and MCD. This underscores the model’s impressive adaptability in handling zero-shot SVS scenarios. 2) StyleSinger 2 also excels in singer similarity, as denoted by the highest MOS-S and Cos. This highlights our model’s superior ability to model and transfer different singing styles precisely, thanks to the innovative design of our components. Our style adaptive decoder effectively improves the rich stylistic details of synthesis quality, rendering the singing voices more natural and of superior quality. Meanwhile, our clustering style encoder shows an excellent capability for modeling styles across a wide range of categories. Finally, the S&D-LM delivers excellent prediction results for style information and phoneme duration, significantly contributing to synthesis quality and singer similarity. As shown in Figure 3, our StyleSinger 2 not only displays greater details in the mel-spectrogram, but also effectively learns the technique, pronunciation, and rhythm of the audio prompt. In contrast, other baseline models lack details in mel-spectrograms, and their pitch curves remain flat, failing to transfer diverse singing styles. Upon listening to demos, it can be found that our model effectively transfers timbre, singing methods, emotion, rhythm, technique, and pronunciation of audio prompts.
@@
Multi-Level Style Control We add global and phoneme-level text embedding to each baseline model to enable style control. Then, we compare StyleSinger 2 using multi-level text prompts. We conduct both parallel and non-parallel experiments according to the target styles. In the parallel experiments, we randomly select unseen audio from the test set, using the GT global style and phoneme-level techniques as the target. In the non-parallel experiments, global styles and six techniques are randomly yet appropriately assigned. For global styles, we specify singing methods (bel canto and pop) and emotions (happy and sad) for each test target. For phoneme-level styles, we select none, one or more specific techniques (mixed voice, falsetto, breathy, vibrato, glissando, and pharyngeal) for each phoneme of target content. As shown in Table 2, we can find that StyleSinger 2 surpasses other baseline models in both the highest synthesis quality (MOS-Q) and style controllability (MOS-C) in both parallel and non-parallel experiments. This indicates that, in addition to excelling in style transfer, our model also performs well in multi-level style control. This success is attributed to our clustering style encoder’s exceptional style modeling capabilities, the S&D-LM’s effective style control, and the style adaptive decoder’s capacity to generate stylistically rich singing voices. Upon listening to demos, it is obvious that our model effectively controls the global singing method and emotion, as well as phoneme-level techniques.
@@
Cross-Lingual Style Transfer To test the zero-shot cross-lingual style transfer performance of various models, we use unseen test data with different lyrics’ languages as prompts and targets for inference (like English and Chinese), using MOS-Q and MOS-S as evaluation. As shown in Table 3, our StyleSinger 2 outperforms other baseline models regarding both synthesis quality (MOS-Q) and singer similarity (MOS-S). Benefiting from our models for comprehensively modeling and effectively transferring diverse styles, StyleSinger 2 performs well in a cross-lingual environment.
Speech-to-Singing Style Transfer We conducted experiments on both parallel and cross-lingual STS style transfer. In parallel experiments, we randomly select samples with unseen singers from the test set as targets and different speech from the same singers to form prompts. In cross-lingual experiments, we select the speech prompt in a different lyric language from the target (such as Chinese and English). As shown in Table 4, we can find that both synthesis quality (MOS-Q) and singer similarity (MOS-S) of StyleSinger 2 are superior to those of baseline models in both parallel and cross-lingual STS experiments. This demonstrates the excellent ability of our model in cross-lingual speech and singing style modeling and transfer.

### v2

# TCSinger: Zero-Shot Singing Voice Synthesis with Style Transfer and Multi-Level Style Control

@@ Abstract
Zero-shot singing voice synthesis (SVS) with style transfer and style control aims to generate high-quality singing voices with unseen timbres and styles (including singing method, emotion, rhythm, technique, and pronunciation) from audio and text prompts. However, the multifaceted nature of singing styles poses a significant challenge for effective modeling, transfer, and control. Furthermore, current SVS models often fail to generate singing voices rich in stylistic nuances for unseen singers. To address these challenges, we introduce TCSinger, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. Specifically, TCSinger proposes three primary modules: 1) the clustering style encoder employs a clustering vector quantization model to stably condense style information into a compact latent space; 2) the Style and Duration Language Model (S&D-LM) concurrently predicts style information and phoneme duration, which benefits both; 3) the style adaptive decoder uses a novel mel-style adaptive normalization method to generate singing voices with enhanced details. Experimental results show that TCSinger outperforms all baseline models in synthesis quality, singer similarity, and style controllability across various tasks, including zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing style transfer. Singing voice samples can be accessed at https://tcsinger.github.io/.
TCSinger: Zero-Shot Singing Voice Synthesis with Style Transfer
@@
Yu Zhang Ziyue Jiang Ruiqi Li Changhao Pan Jinzheng He Rongjie HuangChuxin WangZhou Zhaothanks: Corresponding Author Zhejiang University, Shanghai AI Laboratory {yuzhang34,ziyuejiang,zhaozhou}@zju.edu.cn
@@ 1 Introduction
To address these challenges, we introduce TCSinger, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. TCSinger transfers and controls styles (like singing methods, emotion, rhythm, techniques, and pronunciation) from audio and text prompts to synthesize high-quality singing voices. To model diverse styles (like singing methods, emotion, rhythm, technique, and pronunciation), we propose the clustering style encoder, which uses a clustering vector quantization (CVQ) model to condense style information into a compact latent space, thus facilitating subsequent predictions, as well as enhance both training stability and reconstruction quality. For style transfer and control, we introduce the Style and Duration Language Model (S&D-LM). The S&D-LM incorporates a multi-task language module using audio and text prompts to concurrently predict style information and phoneme duration, thereby enhancing both. To generate singing voices rich in stylistic nuances, we introduce the style adaptive decoder, which employs a novel mel-style adaptive normalization method to refine mel-spectrograms with decoupled style information. Our experimental results show that TCSinger outperforms other current best-performing baseline models in metrics including synthesis quality, singer similarity, and style controllability across various tasks, including zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing (STS) style transfer. Overall, our main contributions can be summarized as follows:
  • We present TCSinger, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. TCSinger excels in personalized and controllable SVS tasks.
@@
  • Experimental results show that TCSinger surpasses baseline models in synthesis quality, singer similarity, and style controllability across various tasks: zero-shot style transfer, multi-level style control, cross-lingual style transfer, and speech-to-singing style transfer.
@@ 2.2 Style Modeling, Transfer and Control
3 TCSinger
Figure 1: The architecture of TCSinger. In Figure (a), S&D-LM represents the Style and Duration Language Model, and LR stands for length regulator. In Figure (b), the S&D-LM autoregressively predicts style information and phoneme duration. In Figure (c), intermediate mel-spectrograms are refined with style information in the style adaptive decoder. In Figure (d), the clustering style encoder extracts style information from mel-spectrograms.
In this section, we first overview the proposed TCSinger. Then, we introduce several critical components, including the clustering style encoder, the style adaptive decoder, and the Style and Duration Language Model (S&D-LM). Finally, we elaborate on the training and inference procedures.
@@ 3.1 Overview
The architecture of TCSinger is depicted in Figure 1(a). We disentangle singing voices into separate representations for content, style (including singing method, emotion, rhythm, technique, and pronunciation), and timbre. For content representation, lyrics are encoded through a phoneme encoder, while a note encoder captures musical notes. For style representation, we use clustering vector quantization (CVQ) in the clustering style encoder to stably condense style information into a compact latent space, thus facilitating subsequent predictions. For timbre representation, we feed a prompt mel-spectrogram, sampled from different audio of the same singer, into the timbre encoder to obtain a one-dimensional timbre vector, disentangling timbre from other information. Then, we utilize the Style and Duration Language Model (S&D-LM) to predict style information and phoneme duration. Since styles and duration of singing voices are closely related, a composite module benefits both. Moreover, the S&D-LM achieves both style transfer and style control with audio and text prompts. Next, we employ the pitch diffusion predictor for F0 prediction and the style adaptive decoder to generate the target mel-spectrogram. The style adaptive decoder generates intricately detailed singing voices using a novel mel-style adaptive normalization method. During training, we train the clustering style encoder for reconstruction in the first phase and S&D-LM for style prediction in the second phase. During inference, we can input audio prompts or text prompts to S&D-LM for style transfer or control. Please refer to Appendix A for more details.
@@ 3.5 Training and Inference Procedures
Training Procedures The final loss terms of TCSinger in the training phase consist of the following parts: 1) CVQ loss $\mathcal{L}_{CVQ}$: the CVQ loss for the clustering style encoder; 2) Pitch reconstruction loss $\mathcal{L}_{gdiff},\mathcal{L}_{mdiff}$: the Gaussian diffusion loss and the multinomial diffusion loss between the predicted and the GT pitch spectrogram for the pitch diffusion predictor; 3) Mel reconstruction loss $\mathcal{L}_{mae},\mathcal{L}_{ssim}$: the MAE loss and the SSIM loss between the predicted and the GT mel-spectrogram for the style adaptive decoder. 4) Duration prediction loss $\mathcal{L}_{dur}$: the MSE loss between the predicted and the GT phoneme-level duration in log scale for S&D-LM in the teacher-forcing mode; 5) Style prediction loss $\mathcal{L}_{style}$: the cross-entropy loss between the predicted and the GT style information for S&D-LM in the teacher-forcing mode.
Figure 2: Inference procedure of TCSinger. In Figure (a), the S&D-LM extracts information from the audio prompt to predict the target style information and phoneme duration, while in Figure (b), the S&D-LM uses multi-level text prompts to predict them.
@@ 4.1 Experimental Setup
In this section, we present the datasets utilized by TCSinger, delve into the implementation and training details, discuss the evaluation methodologies, and introduce the baseline models.
Dataset We use the open-source singing dataset with style annotations, GTSinger (Zhang et al., 2024b), specifically its Chinese and English subset (5 singers, 36 hours of Chinese and English singing and speech). We also enrich our data with M4Singer (Zhang et al., 2022a) (20 singers, 30 hours of Chinese singing), OpenSinger (Huang et al., 2021) (93 singers, 85 hours of Chinese singing), AISHELL-3 (Shi et al., 2021) (218 singers, 85 hours of Chinese speech), and a subset of PopBuTFy (Liu et al., 2022b) (20 singers, 18 hours of English speech and singing). Then, we manually annotate these singing data with global style labels (singing method and emotion) and six phoneme-level singing techniques. Subsequently, we randomly chose 40 singers as the unseen test set to evaluate TCSinger in the zero-shot scenario for all tasks. Notably, our dataset partitioning carefully ensures that both training and test sets for all tasks include cross-lingual singing and speech data. Please refer to Appendix B for more details.
@@
Evaluation Details We use both objective and subjective evaluation metrics to validate the performance of TCSinger. For subjective metrics, we conduct the MOS (mean opinion score) and CMOS (comparative mean opinion score) evaluation. We employ the MOS-Q to judge synthesis quality (including clarity, naturalness, and rich stylistic details), MOS-S to assess singer similarity (in terms of timbre and styles) between the result and prompt, and MOS-C to evaluate style controllability (accuracy and expressiveness of style control). Both these metrics are rated from 1 to 5 and reported with 95% confidence intervals. In the ablation study, we employ CMOS-Q to gauge synthesis quality, and CMOS-S to evaluate singer similarity. For objective metrics, we use Singer Cosine Similarity (Cos) to judge singer similarity, and Mean Cepstral Distortion (MCD) along with F0 Frame Error (FFE) to quantify synthesis quality. Please refer to Appendix C for more details.
@@
Baseline Models We conduct a comprehensive comparative analysis of synthesis quality, style controllability, and singer similarity for TCSinger against several baseline models. Initially, we evaluate our model against the ground truth (GT) and the audio generated by HiFi-GAN (GT (vocoder)). Additionally, we examine TCSinger with two high-performing speech models that conduct style transfer: YourTTS (Casanova et al., 2022) and Mega-TTS (Jiang et al., 2023). To ensure a fair comparison for singing tasks, we enhance these models with a note encoder to process music scores and train them on speech and singing data. Subsequently, we also compare with the best traditional SVS model, RMSSinger (He et al., 2023). Furthermore, we assess TCSinger’s performance against StyleSinger (Zhang et al., 2024a), the first model for zero-shot SVS with style transfer. For more details, please refer to Appendix D.
@@ 4.2 Main Results
Figure 3: Mel-spectrograms depicting the results of zero-shot style transfer. TCSinger effectively captures the rhythm and pronunciation in red boxes, along with the vibrato technique and rhythm in yellow boxes.
@@
Zero-Shot Style Transfer To assess the performance of TCSinger and baseline models in the zero-shot style transfer task, we randomly select samples with unseen singers from the test set as targets and different utterances from the same singers to form prompts. As shown in Table 1, we have the following findings: 1) TCSinger exhibits outstanding synthesis quality, as indicated by the highest MOS-Q and the lowest FFE and MCD. This underscores the model’s impressive adaptability in handling zero-shot SVS scenarios. 2) TCSinger also excels in singer similarity, as denoted by the highest MOS-S and Cos. This highlights our model’s superior ability to model and transfer different singing styles precisely, thanks to the innovative design of our components. Our style adaptive decoder effectively improves the rich stylistic details of synthesis quality, rendering the singing voices more natural and of superior quality. Meanwhile, our clustering style encoder shows an excellent capability for modeling styles across a wide range of categories. Finally, the S&D-LM delivers excellent prediction results for style information and phoneme duration, significantly contributing to synthesis quality and singer similarity. As shown in Figure 3, our TCSinger not only displays greater details in the mel-spectrogram, but also effectively learns the technique, pronunciation, and rhythm of the audio prompt. In contrast, other baseline models lack details in mel-spectrograms, and their pitch curves remain flat, failing to transfer diverse singing styles. Upon listening to demos, it can be found that our model effectively transfers timbre, singing methods, emotion, rhythm, technique, and pronunciation of audio prompts.
@@
Multi-Level Style Control We add global and phoneme-level text embedding to each baseline model to enable style control. Then, we compare TCSinger using multi-level text prompts. We conduct both parallel and non-parallel experiments according to the target styles. In the parallel experiments, we randomly select unseen audio from the test set, using the GT global style and phoneme-level techniques as the target. In the non-parallel experiments, global styles and six techniques are randomly yet appropriately assigned. For global styles, we specify singing methods (bel canto and pop) and emotions (happy and sad) for each test target. For phoneme-level styles, we select none, one or more specific techniques (mixed voice, falsetto, breathy, vibrato, glissando, and pharyngeal) for each phoneme of target content. As shown in Table 2, we can find that TCSinger surpasses other baseline models in both the highest synthesis quality (MOS-Q) and style controllability (MOS-C) in both parallel and non-parallel experiments. This indicates that, in addition to excelling in style transfer, our model also performs well in multi-level style control, and we are the first method for multi-level singing style control. This success is attributed to our clustering style encoder’s exceptional style modeling capabilities, the S&D-LM’s effective style control, and the style adaptive decoder’s capacity to generate stylistically rich singing voices. Upon listening to demos, it is obvious that our model effectively controls the global singing method and emotion, as well as phoneme-level techniques. For more detailed results with objective evaluations, please refer to Appendix E.2.
@@
Cross-Lingual Style Transfer To test the zero-shot cross-lingual style transfer performance of various models, we use unseen test data with different lyrics’ languages as prompts and targets for inference (like English and Chinese), using MOS-Q and MOS-S as evaluation. As shown in Table 3, our TCSinger outperforms other baseline models regarding both synthesis quality (MOS-Q) and singer similarity (MOS-S). Benefiting from our models for comprehensively modeling and effectively transferring diverse styles, TCSinger performs well in a cross-lingual environment.
Speech-to-Singing Style Transfer We conducted experiments on both parallel and cross-lingual STS style transfer. In parallel experiments, we randomly select samples with unseen singers from the test set as targets and different speech from the same singers to form prompts. In cross-lingual experiments, we select the speech prompt in a different lyric language from the target (such as Chinese and English). As shown in Table 4, we can find that both synthesis quality (MOS-Q) and singer similarity (MOS-S) of TCSinger are superior to those of baseline models in both parallel and cross-lingual STS experiments. This demonstrates the excellent ability of our model in cross-lingual speech and singing style modeling and transfer.